OpenStudy (anonymous):

whats the answer to -6 - -22n=3n-(6+5)?

OpenStudy (goldphenoix):

"The answer"... We cannot give you an answer straight-foward, but we can help you. :) We want to simplify a little bit, before we divide. So let's do the parenthesis. What does 3n-(6+5) simplify to?

OpenStudy (anonymous):

18n + 15??

OpenStudy (goldphenoix):

Nope. You did the distributive property wrong. But you can do it another way, an easier way. Since there's no variable in the parenthesis, you can just do 6+5. 6+5 = 11 So it should be 3n - 11, in simplest form. So your equation should now look like: -6 - -22n=3n - 11 Let's do the right side. What does -6 - -22n simplify to? (HINT: Something about the operation)

OpenStudy (goldphenoix):

^If you don't know, I can teach you how. :)

OpenStudy (anonymous):

Would It Be -16?

OpenStudy (goldphenoix):

Nope. You cannot combine those two. Why? Well because one number is a coefficient, which mean a number with a variable. While the other number is not a coefficient. This is why you cannot combine them ^_^ But you can change the operation. a subtraction and a negative becomes a plus sign. :) That's why I said the hint. (I didn't want you to mess up >_>) So your equation should look like: -6 + 22n=3n - 11

OpenStudy (goldphenoix):

Now it's time to combine. What can we combine?

OpenStudy (goldphenoix):

Okay, you cannot combine. My bad. But you know what I mean, right?

OpenStudy (anonymous):

I guess

OpenStudy (goldphenoix):

So how can we simplify it? How can we 'reduce" it, so it can look easier?

OpenStudy (goldphenoix):

Sorry, if I'm not making sense. What I'm trying to say is, " How can we solve for x?"

OpenStudy (goldphenoix):

n*

OpenStudy (anonymous):

We Can subtract from both sides? I think?

OpenStudy (goldphenoix):

Well. We want our result to be n = (A number) So we must bring n to one side. It's easier to bring x to the left so that we won't get a negative. What's 22 divided by 3?

OpenStudy (anonymous):

7

OpenStudy (goldphenoix):

Holy crap! I mest up. I'm so sorry. T~T I thought it was something else. But yes, you are correct. You subtract.

OpenStudy (goldphenoix):

Let restart a little bit. >_> So it looks like: -6 - -22n=3n - 11

OpenStudy (goldphenoix):

You said we can subtract right? So what do we get?

OpenStudy (anonymous):

Oh Ok Then So What Do We Have To Subtract ?

OpenStudy (goldphenoix):

-6 + 22n=3n - 11 We want to find n. So we can subtract the non-coefficient, or the coefficient. Let's do the coefficient first, since I want to. :) What is 22 - 3?

OpenStudy (anonymous):

Its 19

OpenStudy (goldphenoix):

Correct! So our equation should look like: -6 + 19n =- 11 Next, we must subtract -6 and -11. We cannot subtract -11 and -6 because want to bring the -6 to the right side. So what is: -6 - (-11)

OpenStudy (anonymous):

-5 I think?

OpenStudy (goldphenoix):

Nope. Positive 5. Why? Well when you have a minus sign and a negative sign next to eachother, it becomes a positive. So another way of writing it is, -6 + 11. -6 + 11 = 5. (It should be positive because the positive number is greater than the negative number) So your equation should look like: 19n = 5 What do you think we should do?

OpenStudy (anonymous):

Ok Then And We Divide

OpenStudy (goldphenoix):

Ugh. I keep on messing up!!! Yes, you are right. You're suppose to do -11 - 6. That would give you -5. Yes, we do divide! 19n = -5. What is 19n divided by -5?

OpenStudy (goldphenoix):

You deserve a medal, since you literary corrected my mistake. :O

OpenStudy (anonymous):

Thanks And Is It 3.8?

OpenStudy (goldphenoix):

My bad, -5 divided by 19. N = -0.26 I hope this help! (Sorry for wasting like 30 min of your life.)

OpenStudy (goldphenoix):

I know I made a lot of error (You corrected it for me, so the answer is 100% right) So here's a link: http://www.purplemath.com/modules/solvelin3.htm That link explain very well about how to solve these equations. :)

OpenStudy (anonymous):

Its Alright So The Answer Is N= -0.26? And Thanks

OpenStudy (goldphenoix):

Yes. :)

OpenStudy (anonymous):

Ok Thanks So I Just Go On That Website And It helps?

OpenStudy (goldphenoix):

You can leave your answer as \[\large \large N = -\frac{ 5 }{ 19 }\]

OpenStudy (goldphenoix):

Well, it explains. There's no people helping you.

OpenStudy (anonymous):

Ok